stion 15 of 18 (1 point) ese Incorrect iDC Incorrect What is the ratio of acetate ion to acetic acid (K.-1.76 x 10-5) in a buffer containing these compounds at pH 3.570? Number Tools 102 Check Answer:
Answer #1

SOLUTION:

The problem can be solved by Henderson-Hasselbalch equation as:

CH3COOH <--------> CH3COO- + H+

pH = pKa + log [CH3COO-] / [CH3COOH]

pKa = -logKa = - log 1.76 X 10-5 = 4.75; pH = 3.570

3.57 = 4.75 + log[CH3COO-] / [CH3COOH]

log[CH3COO-] / [CH3COOH] = 3.57 - 4.75 = - 1.18

[CH3COO-] / [CH3COOH] = antilog(-1.18) = 0.066

Hence the ratio of acetate to acetic acid is 0.066